### v2

# ICYM2I: The illusion of multimodal informativeness under missingness

Abstract
Multimodal learning is of continued interest in artificial intelligence-based applications, motivated by the potential information gain from combining different types of data. However, modalities observed in the source environment may differ from the modalities observed in the target environment due to multiple factors, including cost, hardware failure, or the perceived informativeness of a given modality. This shift in missingness between the source and target environment has not been carefully studied. Naïve estimation of the information gain associated with including an additional modality without accounting for missingness may result in improper estimates of that modality’s value in the target environment. We formalize the problem of missingness, demonstrate its ubiquity, and show that the subsequent distribution shift results in bias when the missingness process is not explicitly accounted for. To address this issue, we introduce ICYM2I (In Case You Multimodal Missed It), a framework [^1] Available on Github https://github.com/reAIM-Lab/ICYM2I/ for the evaluation of predictive performance and information gain under missingness through inverse probability weighting-based correction. We demonstrate the importance of the proposed adjustment to estimate information gain under missingness on synthetic, semi-synthetic, and real-world datasets.
1 Introduction
Multimodal learning is ubiquitous in machine learning as practitioners combine multiple data types to improve predictive performance in applications to healthcare [1, 2], robotics [3, 4], and recommender systems [5]. However, factors such as privacy concerns [6, 7], cost-benefit tradeoffs of data-acquisition [8], and user preferences [9] imprint multimodal data with missingness. Additionally, even if modality complete data is available or curated at training, data noise [10, 11] and sensor failures [12, 13] may result in missing modalities in the target environment.
Despite the existence of missingness in real-world settings, current multimodal machine learning methods often assume modalities are fully observed, both in source and target environments. When missingness is considered, the literature has focused on engineering efforts [14, 15] such as data selection [16], imputation [17, 18, 19], and architecture design [20, 21], which implicitly assume a stable missingness process between source and target environments. When this assumption is violated, the missingness mechanism induces a distribution shift [22, 23] that biases the estimated informativeness of a given modality. Missingness is pervasive and impacts a broad range of application domains encountered in the multimodal literature: in breast cancer screening, biopsies are only performed if there are abnormal findings in a mammogram; in autonomous vehicles, LiDAR sensor dropout can occur due to weather and lighting conditions; and in online recommender systems, reviews are only collected after certain consumer behaviors. Across these settings, ignoring the distribution shift between source and target due to missingness when quantifying modality informativeness may conflate missingness with signal, leading to flawed data collection and modeling decisions.
In this work, we propose a framework to overcome the (mis)estimation of both inherent informativeness and predictive utility under missingness in multimodal learning. Our contributions, summarized in Figure 1, are as follows:
  • Framework for multimodal learning with missingness. We formalize the impact of missingness as a distribution shift intrinsic to multimodal learning, where the observed source distribution differs from the target distribution due to missingness. We show that not accounting for missingness, a common practice, may bias the estimate of a modality’s predictive and information-theoretic utility.
  • ICYM2I. Under the missingness-at-random (MAR) assumption, a much more realistic assumption than the common and often implicit assumption of missingness-completely-at-random (MCAR) made by state-of-the-art multimodal strategies, we propose ICYM2I  (In Case You Multimodal Missed It), a double inverse-propensity weighting correction to overcome missingness-induced distribution shifts. Specifically, we demonstrate that ICYM2I improves correlation in predictive and information-theoretic utility of modalities.
  • Experiments on diverse data. We demonstrate the broad applicability and utility of our methods in synthetic, semi-synthetic, and real-world benchmark datasets, including a case study in multimodal learning in health.
Figure 1: Overview of the proposed framework. Curation often discards missing data, resulting in a discrepancy between the collected $\Omega$ and source datasets $\Omega_{\text{source}}$ used for training. Current practice is denoted in blue: naïve training and evaluating on $\Omega_{\text{source}}$ leads to biased estimates of performance and informativeness on target data. The orange path illustrates the proposed ICYM2I: a double inverse probability weighting (IPW) mechanism that yields accurate performance and informativeness estimates under the target distribution.
2 Related Work
Multimodal benchmarks suppress missingness encountered in real-world environments. Prior work on multimodal models often assumes fully observed modalities [24, 25, 26]. Missingness has largely been an overlooked problem [14, 15], to the extent that current benchmarks rarely contain samples with missing modalities. Curation often involves dropping incomplete or filtering samples based on data quality criteria, such as text length or file size [27, 28] or imputing with automatic tools [29]. This curation implicitly assumes that rejected samples follow the same distribution as the observed ones. This assumption may not hold. For example, in autonomous driving data, samples with sensor failure – often resulting from extreme weather or lighting conditions – may be filtered out. Models trained on complete data may, consequently, not generalize to these settings, creating real-world risk at deployment. When missingness is considered, previous works focused on robustness through imputation [17, 18, 19], representation learning [15, 30], knowledge distillation [31, 32], and model ensembling [20, 21] – all ignoring the potential shift resulting from the missingness process.
Multimodal missingness in the target environment. Prior work has explored missingness in the target environment [33, 21], e.g., when a captor fails at deployment [34]. Broadly, two strategies have been proposed [15]: (i) data preprocessing through cross-modal imputation [18, 19, 17], where one replaces the missing modality [35, 36], as well as (ii) model training strategies such as architecture design [37, 38], distillation-based methods [31, 32], and ensembling [20, 21]. Through the proposed formalization, our work distinguishes between different missingness assumptions, demonstrating that the previously studied framework is only one among various plausible mechanisms for which current strategies are not well-designed.
Distribution shifts in multimodal learning. Addressing multimodal shifts has been studied in vision-language models [39, 40] or using information-theoretic notions to understand multimodal behavior under distribution shifts [41]. Augmentation and regularization strategies have been leveraged to address temporal shifts for conversation understanding [42, 43]. Advances in learning, such as in-context learning, have been studied to characterize adaptation to multimodal distribution shifts [39, 44]. However, existing strategies aim to improve robustness under domain distribution shifts, while ignoring the potential shift in missingness between source and target environments.
Quantifying information-theoretic value of a modality. Existing works often implicitly assume that additional modalities improve performance, ignoring the prohibitive cost, complexity, and potential noise added by these additional dimensions. When limited resources or constraints limit availability in the target environment, a central challenge is to quantify the information-theoretic value of a modality [45]. Liang et al. [46] proposed a method for recovering partial information decomposition measures of the redundancy, uniqueness, and synergy of the information provided by the different modalities [47, 48]. However, these decompositions fully ignore the impact of missingness.
Correcting for missingness bias. The lack of formalization of missingness in the multimodal literature has led to neglecting its potential impact. Ignoring this process risks biasing estimates of interest [49] as the observed distribution differs from the underlying one practitioners aim to model. The statistical literature has introduced strategies such as matching [50] and reweighting [51] to correct for the missingness process. However, these works have overlooked the multimodal setting and the systematic shifts that may occur in this setting.
3 Multimodality and missingness
[…]
Figure 2: Directed Acyclic Graphs of the assumed data-generating processes. On the left is the commonly assumed graph with no missingness. On the right is the proposed missingness formalism. $X_{1}$ and $X_{2}$ are two modalities of interest, $Y$ is the label of interest. The missingness process depends on $C$. Filled point nodes are observed variables, while unfilled nodes are unobserved. Gray edges indicate MAR missingness for a given modality.
Consider two modalities, $X_{1}\in\mathcal{X}_{1}$ and $X_{2}\in\mathcal{X}_{2}$ and the state of interest $Y\in\mathcal{Y}$. We denote the joint underlying distribution $\Omega=\mathcal{X}_{1}\times\mathcal{X}_{2}\times\mathcal{Y}$. Without loss of generality, we assume an anti-causal setting for the data-generating process, in which the modalities are dependent on the states $Y$, as shown in Figure 2 (left). We use the binary indicators of missingness $M_{1}$, $M_{2}$, and $M_{Y}$, which are equal to $1$ if the associated variable is missing, 0 if observed, following the convention of Mohan and Pearl [52]. Observed variables are subscripted by ‘obs’, which corresponds to the underlying modality if observed, and unobserved otherwise (denoted by $\varnothing$). Formally, the observed variable $Y_{obs}$ and observed modalities $X_{1,obs}$ and $X_{2,obs}$ can be defined as follows:
$Y_{obs}=\begin{cases}\varnothing&\text{ if }M_{Y}=1,\\ Y&\text{ otherwise.}\end{cases}$
In this setting, we denote the observed joint distribution $\Omega_{\text{obs}}^{\text{}}:=(M_{1}\cdot\mathcal{X}_{1})\times(M_{2}\cdot\mathcal{X}_{2})\times(M_{Y}\cdot\mathcal{Y})$, which has been the focus of multimodal learning. Note that the underlying relation between the covariates of interest $X_{1}$, $X_{2}$, and $Y$ remains unchanged under the true data-generating mechanism; only the realizations of the observed variables differ.
Missingness in multimodal learning. We distinguish three mechanisms that cover potential missingness in multimodal data [53]:
  • Missing Completely At Random (MCAR): A modality is missing completely at random if the missingness process is independent of any other variable.
  • Missing At Random (MAR): The missingness mechanism depends on observed variables only.
  • Missing Not At Random (MNAR): Missingness depends on unobserved variables.
In Figure 2 (right), we describe the missingness mechanisms as dependent on $C$, a set of covariates that determine the missingness mechanism. Note that $C$ may include one of the modalities of interest, e.g., whether $X_{2}$ is observed may depend on the realization of $X_{1}$. In general, the set $C$ may differ for each modality depending on the data-generating process.
Missingness-induced distribution shifts. Missingness in a modality $X_{i}$ and/or the label $Y$ can induce distribution shifts between the source and the target distributions. For example, if a modality is observed in the target environment only if another one meets some criterion, then this distribution may not match the source distribution. Theoretically, we know that a non-MCAR missingness mechanism induces distribution shifts [23, 22], i.e., the observed distribution differs from the underlying distribution. Critically, models trained and evaluated on the observed distribution are statistically biased estimates under any other missingness process.
For instance, consider an autonomous vehicle setting where video and LiDAR represent two modalities of interest. If the LiDAR randomly dysfunctions, the missingness patterns are MCAR. However, as previously mentioned, LiDAR may malfunction under extreme weather conditions. If these conditions can be extracted from the video modality, one may assume MAR patterns. However, if the video cannot capture the variable explaining the LiDAR dysfunction – e.g., temperature – LiDAR would be MNAR, as dependent upon an unobserved variable. Under the last two scenarios, focusing solely on samples with LiDAR excludes all extreme condition settings.
A common and often implicit assumption in the multimodal literature is the absence of missingness, which is analogous to either a MCAR mechanism or a stable missingness process as described in Assumption A. In other words, not adjusting for missingness assumes that the missingness process is uninformative or will remain the same in the target environment.
Assumption A (Stable missingness process).
The missingness process is stable between the source and target environments, i.e., the observed distributions are the same $\Omega_{\text{obs}}^{\text{source}}=\Omega_{\text{obs}}^{\text{target}}$.
A few prior works focus on improving the robustness of multimodal models when performance may degrade due to a modality missing in the target distribution, as formalized in Assumption B.
Assumption B (Collection degradation).
Collection in the target environment may deteriorate, resulting in a distribution shift $\Omega_{\text{obs}}^{\text{source}}\sim\Omega^{\text{}}\neq\Omega_{\text{obs}}^{\text{target}}$.
Our work questions the applicability of these assumptions where modality collection is costly. While missingness may naturally present a distributional shift [54], we emphasize that demonstrating the value of a modality in the source environment may lead to increased collection of this modality in the target environment, inducing a distribution shift akin to Assumption C.
Assumption C (Multimodal analysis informs data collection).
Demonstrated multimodal performance gain induces a shift in the missingness process in the target, i.e. $\Omega_{\text{obs}}^{\text{source}}\neq\Omega^{\text{target}}\sim\Omega^{\text{}}$.
We focus on settings where historical data on which one can train a model is marked by missingness. Under such settings, we aim to do the following: (i) identify which modalities are informative and may consequently be fully observed in the target environment, and (ii) train models that generalize to the full observation of these modalities.
4 Is this modality informative?
We aim to assess whether a partially missing modality would be informative if fully observed. To this end, we introduce ICYM2I (In Case You Multimodal Missed It), a framework for correcting model performance trained on modality complete samples where all modalities and labels are observed $\left(\Omega_{\text{obs}}^{\text{}}\right)$ to estimate the predictive utility of the partially missing modality if it were observed for the whole population ($\Omega^{\text{}}$). Additionally, we propose a correction to derive unbiased estimates of the information-theoretic utility of a modality, using $\Omega_{\text{obs}}^{\text{}}$. We rely on Partial Information Decomposition (PID) [48] bounds introduced by Bertschinger et al. [47] for this task, which quantifies the information value of a target of interest captured by two input variables.
Correction. We propose an Inverse Probability Weighting (IPW) variant [55], which reweights samples based on their probability of being observed. Under Assumption D, IPW recovers unbiased estimates of the true distribution, enabling learning and evaluation on the true distribution from observed samples. IPW-adjustment is critical for both training and evaluation of multimodal models under missingness. IPW-adjusted training results in a model trained to infer on the underlying distribution $\Omega^{\text{}}$, while correction of the evaluation allows for measuring performance on $\Omega^{\text{}}$, despite evaluating the model only on samples from the observed distribution $\Omega_{\text{obs}}^{\text{}}$.
Assumption D (MAR and Positivity).
The missingness mechanism is MAR, and $p_{\Omega^{\text{}}}(M_{1}=0,M_{2}=0,M_{Y}=0|C)>0$.
4.1 A motivating example
We consider the common multimodal example of learning bit-wise logic operators [47, 56, 46]. We generate $10,000$ points with two modalities drawn from Bernoulli distributions ($p=0.5$). The output state $Y$ is defined using the binary operators AND, OR, and XOR of input bits $X_{1}$ and $X_{2}$. In this setting, we induce missingness $M_{2}$ in $X_{2}$ and $Y$ as a function of $X_{1}$ (MAR): $M_{2}\sim Bern(0.6X_{1}+0.2)$, resulting in 50% missingness in $X_{2}$. We investigate the impact of missingness on current strategies for evaluating the predictive and information-theoretic utility of a given modality.
Estimating performance for informativeness. A common practice to measure the predictive value of adding a modality is through modality ablation studies where practitioners train models on the subset of observed samples where all modalities are observed $\left(\Omega_{\text{obs}}^{\text{}}\right)$. First, unimodal models $f_{\phi}(x_{i})$ are trained to approximate $p_{\Omega_{\text{obs}}^{\text{}}}(y\mid x_{i}),\,\forall i\in\{1,2\}$ and a multimodal model $f_{\phi}(x_{1},x_{2})$ to approximate $p_{\Omega_{\text{obs}}^{\text{}}}(y\mid x_{1},x_{2})$ on the same observed dataset. The performance is then compared in a hold-out set, which is also sampled from $\Omega_{\text{obs}}^{\text{}}$. The informativeness of a modality is heuristically attributed to the relative performance gain of the multimodal model compared to the unimodal model. However, multimodal models can perform worse than their unimodal counterparts due to data characteristics [57] and learning dynamics [58, 59]. Thus, relying solely on performance as a proxy for informativeness, particularly under distribution shifts, can be misleading.
Partial Information Decomposition [47]. As an alternative to estimating performance, existing works have decomposed the informativeness associated with each modality [46]. Partial information decomposition (PID) [47, 48] decomposes the total mutual information $I(Y:(X_{1},X_{2}))$ [60, 61] between a random variable $Y$ and $(X_{1}$, $X_{2})$ jointly into shared information (information both $X_{1}$, $X_{2}$ share about $Y$), unique information 1 (information only $X_{1}$ has about $Y$), unique information 2 (information only $X_{2}$ has about $Y$), and complementary information (information about $Y$ that requires both $X_{1}$ and $X_{2}$):
$I(Y:(X_{1},X_{2}))=$ $\underbrace{SI(Y:X_{1};X_{2})}_{\text{shared information}}+\underbrace{UI(Y:X_{1}\backslash{X_{2}})}_{\text{unique information 1}}+\underbrace{UI(Y:X_{2}\backslash{X_{1}})}_{\text{unique information 2}}+\underbrace{CI(Y:X_{1};X_{2})}_{\text{complementary information}}$
Let $\Delta$ be the space of all distributions over $(X_{1},X_{2},Y)$ and let $\Omega^{\text{}}$ denote the true data distribution (without missingness) and define $\Delta_{\Omega^{\text{}}}:=\big\{q\in{\Delta}:q(X_{i}=x_{i},Y=y)=p_{\Omega}^{\text{}}(X_{i}=x_{i},Y=y)\text{ }\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}\big\}$. That is, $\Delta_{\Omega^{\text{}}}$ is the set of all distributions over $(X_{1},X_{2},Y)$ such that the two-way joints between $X_{i}$ and $Y$ match the true data-generating distribution. Equipped with this set, Bertschinger et al. [47] provides the following bounds $\widetilde{SI}$, $\widetilde{UI}$, and $\widetilde{CI}$ on the analogous quantities:
$\widetilde{SI}(Y:X_{1};X_{2})$ $:=\max_{Q\in{\Delta_{\Omega^{\text{}}}}}[I_{q}(Y:X_{1})-I_{q}(Y:X_{1}|X_{2})]$ $\leq SI(Y:X_{1};X_{2})$ (1)
[…]
$\widetilde{CI}(Y:X_{1};X_{2})$ $:=I_{\Omega^{\text{}}}(Y:(X_{1},X_{2}))-\min_{Q\in{\Delta_{\Omega^{\text{}}}}}I_{q}(Y:(X_{1},X_{2}))$ $\leq CI(Y:X_{1};X_{2})$ (4)
These bounds are tight if there exists a $q_{0}\in\Delta_{\Omega^{\text{}}}$ such that $\widetilde{CI}_{q_{0}}(Y:X_{1};X_{2})=0$. Bertschinger et al. [47] further shows that under common (but unverifiable) assumptions on the data-generating process, the inequalities are tight for all $q\in\Delta_{\Omega^{\text{}}}$. This results in a compelling argument to rely on these entities, as it suggests that it is not possible to decide whether or not complementary information exists when only marginals $(Y,X_{1})$ and $(Y,X_{2})$ are known. Hence, we choose this measure of PID. Prior works relying on this PID to attribute information-theoretic value implicitly assume that $\Omega_{\text{obs}}^{\text{source}}=\Omega^{\text{source}}=\Omega^{\text{target}}=\Omega$. Instead, we evidence the limitations of these strategies performed on $\Omega_{\text{obs}}^{\text{source}}$ when the target decomposition is $\Omega^{\text{target}}=\Omega$, i.e., the true data-generating mechanism.
Table 1: Impact of missingness on multimodality information for bitwise logic operators. Parentheses denote standard deviation across batches.
AUROC Information Decomposition
$X_{1}$ $X_{2}$ $X_{1}+X_{2}$ Unique 1 Unique 2 Shared Complementary
AND Oracle 0.83 (0.01) 0.84 (0.01) 1.00 (0.00) 0.05 (0.00) 0.03 (0.00) 0.26 (0.00) 0.47 (0.00)
Observed 0.66 (0.01) 0.93 (0.01) 1.00 (0.00) 0.44 (0.00) 0.00 (0.00) 0.15 (0.00) 0.36 (0.00)
ICYM2I 0.83 (0.01) 0.85 (0.02) 1.00 (0.00) 0.03 (0.00) 0.03 (0.00) 0.27 (0.00) 0.45 (0.00)
OR Oracle 0.84 (0.01) 0.83 (0.01) 1.00 (0.00) 0.04 (0.00) 0.05 (0.00) 0.27 (0.00) 0.46 (0.00)
Observed 0.95 (0.01) 0.77 (0.01) 1.00 (0.00) 0.01 (0.00) 0.15 (0.00) 0.10 (0.00) 0.23 (0.00)
ICYM2I 0.85 (0.02) 0.82 (0.01) 1.00 (0.00) 0.03 (0.00) 0.02 (0.00) 0.27 (0.00) 0.50 (0.00)
XOR Oracle 0.51 (0.02) 0.49 (0.01) 1.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.99 (0.00)
Observed 0.52 (0.02) 0.80 (0.02) 1.00 (0.00) 0.34 (0.00) 0.07 (0.00) -0.07 (0.00) 0.62 (0.00)
ICYM2I 0.53 (0.03) 0.49 (0.03) 1.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.01 (0.00) 0.96 (0.00)
As a motivating example, we analyze the impact of missingness on estimating PID in the case of unidimensional modalities with a bitwise logic outcome (AND, OR, and XOR). Table 1 (left) presents the discriminative performance associated with neural networks trained on each modality and their combination under three scenarios: (i) access to all data (Oracle), (ii) focusing only on datapoints with all covariates observed (Observed), and (iii) adequately accounting for missingness (ICYM2I  using IPW to adjust $\Omega_{\text{obs}}^{\text{}}\mapsto\Omega^{\text{}}$, by modeling the missingness mechanism), as proposed in Section 4.2. Table 1 (right) presents PID, discussed in Section 4.3 under the same scenarios, demonstrating how information decomposition is also biased due to missingness.
Specifically, relying on $\Omega_{\text{obs}}^{\text{}}$ overestimates the performance of $X_{1}$ for OR but underestimates it for AND. Similarly, biased decomposition results in overestimating the informativeness of $X_{1}$ (“Unique 1” compared to “Unique 2”) for OR. As $X_{1}$ informs the missingness process, it indirectly informs the outcome of interest, despite the true underlying generative process being dependent on both. The use of IPW can correct for such bias under positivity as long as the propensities for IPW can be estimated (i.e., the MAR assumption). We study sensitivity to this assumption in Appendix D, where we further evaluate the robustness of our method under MCAR and MNAR, demonstrating robustness under MCAR.
We now formally describe two methods for reliably inferring the informativeness of modalities using (i) unbiased estimation of unimodal versus multimodal model performance using supervised learning (ICYM2I-learn), and (ii) high-dimensional autodifferentiable partial information decomposition (ICYM2I-PID). In addition, we demonstrate the need for IPW-adjusted evaluation as a key element to determine modality informativeness using supervised learning.
4.2 ICYM2I-learn: Estimating performances for informativeness under missingness
Training. Under the MAR assumption, i.e., the missingness is fully explained by observed covariates $C$; that is, the probability of a data point being missing depends only on $C$, we propose to train the model with a weighted loss using samples from $\Omega_{\text{obs}}^{\text{}}$. The proposed IPW-adjusted loss accounts for the distributional shift ($\Omega_{\text{obs}}^{\text{}}\mapsto\Omega^{\text{}}$) by up-weighting under-observed points, as described in the following lemma.
Lemma 1 (IPW Training).
The loss function computed on the observed data $l_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2},y)$ can be reweighted to approximate the target loss $l_{\Omega}^{\text{}}(x_{1},x_{2},y)$ as follows:
$l_{\Omega}^{\text{}}(x_{1},x_{2},y)=\frac{1}{1-p(m_{1},m_{2},m_{y}\mid C)}\,l_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2},y)$
where $p(m_{1},m_{2},m_{y}\mid C)$ is the probability of missingness, given the covariates $C$.
Evaluation. Current works suffer from an analogous bias in model evaluation, by relying on a hold-out set from the observed distribution ($\Omega_{\text{obs}}^{\text{}}$). To estimate a given metric on the true underlying distribution, one must correct this metric using a similar correction as previously described. Li et al. describes how to correct for both AUC and Brier score using IPW.
Corollary 1 (ICYM2I-learn).
Consider a model $f$ trained and evaluated on data drawn from $\Omega_{\text{obs}}^{\text{}}$. To correct the model and estimate its performance on $\Omega^{\text{}}$, one must correct both its training and evaluation following the previous corrections.
4.3 ICYM2I-PID: Partial information decomposition for multimodal informativeness
Under missingness, we have samples from $\Omega_{\text{obs}}^{\text{}}$ instead of $\Omega^{\text{}}$. Estimating PID measures in this setting requires adjusting for the $\Omega_{\text{obs}}^{\text{}}\mapsto\Omega^{\text{}}$ shift. The three-way mutual information is the key estimand of interest to obtain any PID bound, requiring the following correction (derived in Appendix A):
Lemma 2 (Corrected mutual information).
$I_{\Omega^{\text{}}}(Y:(X_{1},X_{2}))=\mathbb{E}_{\begin{subarray}{c}x_{1},x_{2}\sim p_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2})\\ y\sim p_{\Omega^{\text{}}}(y\mid x_{1},x_{2})\end{subarray}}\left[\frac{1-p(m_{1},m_{2})}{1-p(m_{1},m_{2}|x_{1},x_{2},y)}\log{}\left(\frac{p_{\Omega^{\text{}}}(x_{1},x_{2},y)}{p_{\Omega^{\text{}}}(x_{1},x_{2})p_{\Omega^{\text{}}}(y)}\right)\right]$
and analogously for other quantities. Second, these mutual information-based quantities can be equivalently estimated using entropy-based measures (see derivations in Appendix B). Third, to enable PID for high-dimensional data, we parameterize $q$ via unimodal neural networks $f_{1}$ and $f_{2}$, motivated by Liang et al. [46]. Then,
$\Delta_{\Omega^{\text{}}}\approx\{q\propto\exp(f_{1}(x_{1})\cdot f_{2}(x_{2})):q(x_{i},y)={p_{\Omega^{\text{}}}}_{\phi}(x_{i},y)\,\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}\}$
in which $\Omega^{\text{}}_{\phi}$ is a re-parametrization of $\Omega^{\text{}}$ using neural networks and learned using samples from $\Omega_{\text{obs}}^{\text{}}$. That is, prior work considers ${p_{\Omega^{\text{}}}}_{\phi}(y,x_{i})\approx f_{\phi}(y\mid x_{i})p_{\Omega_{\text{obs}}^{\text{}}}(x_{i}),\forall i\in\{1,2\}$, with $f_{\phi}$, a unimodal neural network, trained on samples from $\Omega_{\text{obs}}^{\text{}}$ to predict $Y$.
Due to missingness, ${\Omega^{\text{}}}_{\phi}\neq{\Omega_{\text{obs}}^{\text{}}}_{,\phi}$, and should be approximated using IPW corrections for the ${\Omega^{\text{}}}_{\phi}\mapsto{\Omega_{\text{obs}}^{\text{}}}_{,\phi}$. We use the modified Sinkhorn-Knopp algorithm [63] to enforce matching the two-way IPW-corrected marginals to constraint $q\in\Delta_{\Omega^{\text{}}}^{\textsc{ICYM${}^{2}$I}}$ [^2] Prior works match conditionals using samples from $\Omega_{\text{obs}}^{\text{}}$ without correction: $q(y\mid x_{i})\approx\Omega_{\text{obs}}^{\text{}}(y\mid x_{i})$ which may produce biased PID bounds.. Using the fact that all PID bounds can be obtained by minimizing the three-way mutual information (see Bertschinger et al. [47], Lemma 4), we summarize our algorithm to minimize $I_{q}(Y:(X_{1},X_{2}))$ in Appendix C.
5 Experiments
To better understand the connection between performance, information decomposition and missingness, we propose a simulation (detailed in Appendix E), two semi-synthetic studies that reflect real-world missingness mechanisms (see Appendix F), and a real-world case-study.
5.1 Simulation and semi-synthetic experiments
In Table 2, each point represents the estimated PID value (Unique 1, Unique 2, Shared, and Complementary) for one simulation under the training and evaluation IPW-corrections and the oracle performance, i.e., a model trained and tested on $\Omega^{\text{}}$. Specifically, columns reflect evaluation correction, while rows reflect training correction. These results underline the importance of correcting both training and evaluation, as proposed in ICYM2I, to best align with the performance one would obtain on $\Omega^{\text{}}$, as shown by the smallest Root Mean Squared Error (RMSE) observed when both corrections are applied. This observation shows that the proposed ICYM2I  best recovers the true informativeness of each modality, despite relying on $\Omega_{\text{obs}}^{\text{}}$. Appendix E echoes the same observation when evaluating model performance.
Table 2: Comparison between estimated PID using training and PID corrections, and oracle PID on $\Omega^{\text{}}$. $\epsilon$ denotes the RMSE between estimated and oracle PIDs.
Observed ($\Omega_{\text{obs}}^{\text{}}$) Underlying ($\Omega^{\text{}}$)
$\Delta_{\Omega}^{\text{}}$-Correction $\Omega_{\text{obs}}^{\text{}}$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$\Omega^{\text{}}$ [Uncaptioned image] [Uncaptioned image]
The semi-synthetic experiments examine the effect of enforcing increasing missingness on the performance and information decomposition of UR-FUNNY [64] and hateful memes [65], two foundational real-world datasets used in the multimodal literature for affective computing and content moderation. Table 3 summarizes the effect of enforcing 70% missingness on estimating multimodality informativeness across these datasets, demonstrating the generalizability of our proposed strategy across real-world datasets. Appendix F further illustrates the robustness of the methodology under different levels of missingness in these datasets.
Table 3: Impact of 70% missingness on multimodality information for UR-FUNNY [64] and Hateful Memes [65]. Parentheses denote standard deviation across batches.
AUROC Information Decomposition
Text Image/Video Image + Text Unique Text Unique Image Shared Complementary
UR-FU. Oracle 0.68 (0.01) 0.60 (0.02) 0.69 (0.02) 0.10 (0.00) 0.02 (0.00) 0.00 (0.00) 0.00 (0.00)
Observed 0.61 (0.03) 0.54 (0.04) 0.63 (0.03) 0.05 (0.00) 0.00 (0.00) 0.03 (0.00) 0.00 (0.00)
ICYM2I 0.66 (0.03) 0.57 (0.04) 0.62 (0.04) 0.07 (0.00) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00)
Memes Oracle 0.71 (0.01) 0.57 (0.01) 0.72 (0.01) 0.09 (0.01) 0.00 (0.00) 0.04 (0.00) 0.05 (0.01)
Observed 0.68 (0.02) 0.61 (0.02) 0.71 (0.02) 0.13 (0.00) 0.04 (0.00) 0.01 (0.00) 0.00 (0.00)
ICYM2I 0.67 (0.02) 0.61 (0.02) 0.71 (0.02) 0.10 (0.00) 0.01 (0.00) 0.02 (0.03) 0.03 (0.01)
5.2 Chest radiographs are uninformative over electrocardiograms for structural heart disease detection.
Structural heart disease (SHD) is a set of conditions that affect the heart’s physiology and is typically diagnosed using transthoracic echocardiograms (TTEs) [66]. However, TTEs are often underutilized in the United States due to diagnostic stewardship and competing financial incentives [67]. Prior work using unimodal models with common modalities in electrocardiograms (ECGs) [68, 69] and chest radiographs (CXRs) [70] has demonstrated that non-TTE modalities can detect structural heart disease labels. However, CXRs are not systematically collected in conjunction with ECGs, leading to systematic missingness patterns. We, therefore, evaluate ICYM2I on this clinical task to evaluate the informativeness of CXRs in diagnosing SHD, despite its missingness.
Dataset. Our study population consists of a retrospective study gathering 98,397 adult patients who received an ECG and a TTE within one year of each other. The population has 20.56% SHD prevalence. In this cohort, 12,587 members (12.79%) have recorded CXRs. For subjects with multiple echocardiograms, we select the first TTE to model opportunistic screening with non-TTE modalities. All data were collected from an academic urban medical system between 2008 and 2022. We generate embeddings using modality-specific foundation models–ECG embeddings are generated using ECG-FM [71] and CXR embeddings with ELIXR-C [72]. Data are split temporally, where subjects with TTEs collected on or after 2018 ($n=40,734$) are allocated to the test set. All data were de-identified, retrospective, and collected for clinical purposes from an academic hospital system, with approval from the Institutional Review Board. Appendix G contains further details regarding preprocessing and embedding generation.
Results. Table 4 presents the performance of each uni- and multimodal model, along with the associated information decomposition. While both the observed and corrected analyses demonstrate the importance of ECG in modeling SHD, the corrected results raise questions about the information gain associated with CXR. Naive decomposition suggests the unique information in CXRs at about 5% of the total information. However, ICYM2I  reduces this unique contribution to 1.8% while increasing estimates of shared information between ECG and CXRs for SHD detection. In contrast to domain knowledge, where ECGs capture electrophysiology while CXRs capture structure and anatomy, two distinct aspects of cardiac health, the corrected complementary and shared results, and low unique information of CXRs suggest that CXRs are not independently useful for SHD diagnosis. Note that our results indicate that the multimodal model performs slightly worse than the unimodal ECG model, reflecting the potential overfitting risk associated with a large number of features.
Table 4: Informativeness of ECG and CXR modalities on model-based structural heart disease detection. Parentheses denote standard deviation across batches ($n=1024$).
AUROC Information Decomposition
ECG CXR ECG + CXR Unique ECG Unique CXR Shared Complementary
Observed 0.83 (0.01) 0.72 (0.02) 0.82 (0.01) 0.11 (0.00) 0.01 (0.00) 0.10 (0.00) -0.00 (0.00)
ICYM2I 0.82 (0.01) 0.73 (0.02) 0.83 (0.01) 0.07 (0.00) 0.01 (0.00) 0.48 (0.00) 0.01 (0.00)
6 Discussion
This work formalizes the issue of missingness when considering multiple modalities. We emphasize that existing works commonly overlook modality missingness by discarding samples with missingness at the curation stage, or implicitly assume that the missingness mechanism remains stable when a model is deployed in the target environment. Our work formalizes this problem and demonstrates its ubiquity in the multimodality literature. Most critically, prior work ignores the fact that any perceived informativeness of a modality may result in increased rates of data collection, inducing different missingness patterns at deployment. Our work, therefore, introduces ICYM2I, a correction to estimate the information gain associated with a partially observed modality. Our results demonstrate the methodology’s capacity to correct for biases introduced by missingness across synthetic, semi-synthetic, and real-world multimodal datasets. Finally, we evidence the practical utility of this methodology in a healthcare dataset, demonstrating the diverging conclusions that one would reach if ignoring missingness. Our work highlights the critical importance of missingness in multimodal research and urges practitioners to pay particular attention to this issue by systematically collecting data with incomplete modalities and carefully modeling and accounting for missingness to enhance robustness.
Limitations. The key assumption in our work is that modalities are MAR. Practitioners should ensure that this assumption is appropriate for their data. Importantly, MAR is less restrictive than the implicit MCAR assumption made in the multimodal literature. Furthermore, accounting for MNAR requires distributional assumptions that are rarely met or even testable in real-world settings. Additionally, our work is based on Partial Information Decomposition (PID [47]), which focuses on two input modalities. While theoretical advances to extend to multiple modalities remain an open question [73, 74], in practice, practitioners should consider a one-vs-all approach to inform modality informativeness using our method. Like in prior work on PID-based measures of information gain on high-dimensional data [75, 46], the quality of the representations used may impact the measures returned by ICYM2I. We ensure that our probabilistic estimates are calibrated to mitigate any such challenges.
Acknowledgments
VJ, YSC, and SJ would like to acknowledge partial support from NIH R01MH137679. SJ would like to acknowledge partial support from the SNF Center for Precision Psychiatry and Mental Health at Columbia. Any opinions, findings, conclusions, or recommendations in this manuscript are those of the authors and do not reflect the views, policies, endorsements, expressed or implied, of any aforementioned funding agencies/institutions. The study was approved under IRB-AAAU7973 at Columbia University.
Appendix A Proofs
This section provides the proofs for Lemma 1 and Lemma 2.
Lemma 1.
The separable loss function computed on the observed data $l_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2},y)$ can be reweighted to approximate the target loss $l_{\Omega}^{\text{}}(x_{1},x_{2},y)$ as follows:
$l_{\Omega}^{\text{}}(x_{1},x_{2},y)=\frac{1}{1-p_{\Omega^{\text{}}}(m_{1},m_{2},m_{y}\mid C)}\,l_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2},y)$
where $p_{\Omega^{\text{}}}(m_{1},m_{2},m_{y}\mid C)$ is the probability of missingness, given the covariates $C$.
Proof.
The proof is analogous to that of Lemma 2, which we show in detail, for any separable loss function $l(x_{1},x_{2},y)$. ∎
Lemma 2.
$I_{\Omega^{\text{}}}(Y:(X_{1},X_{2}))=\mathbb{E}_{\begin{subarray}{c}x_{1},x_{2}\sim p_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2})\\ y\sim p_{\Omega^{\text{}}}(y\mid x_{1},x_{2})\end{subarray}}\left[\frac{1-p(m_{1},m_{2})}{1-p(m_{1},m_{2}|x_{1},x_{2},y)}\log{}\left(\frac{p_{\Omega^{\text{}}}(x_{1},x_{2},y)}{p_{\Omega^{\text{}}}(x_{1},x_{2})p_{\Omega^{\text{}}}(y)}\right)\right]$
Proof.
Let $m=(m_{1},m_{2})$,
$I_{\Omega^{\text{}}}(Y:(X_{1},X_{2}))$
$=\mathbb{E}_{\Omega^{\text{}}}\left[\log{}\left(\frac{p_{\Omega}^{\text{}}(x_{1},x_{2},y)}{p_{\Omega}^{\text{}}(x_{1},x_{2})p_{\Omega}^{\text{}}(y)}\right)\right]$
$=\mathbb{E}_{\begin{subarray}{c}x_{1},x_{2}\sim p_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2})\\ y\sim p_{\Omega^{\text{}}}(y\mid x_{1},x_{2})\end{subarray}}\left[\frac{p_{\Omega^{\text{}}}(x_{1},x_{2},y)}{p_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2},y)}\log{}\left(\frac{p_{\Omega}^{\text{}}(x_{1},x_{2},y)}{p_{\Omega}^{\text{}}(x_{1},x_{2})p_{\Omega}^{\text{}}(y)}\right)\right]$
$=\mathbb{E}_{\begin{subarray}{c}x_{1},x_{2}\sim p_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2})\\ y\sim p_{\Omega^{\text{}}}(y\mid x_{1},x_{2})\end{subarray}}\left[\frac{p_{\Omega}^{\text{}}(x_{1},x_{2},y)}{p_{\Omega}^{\text{}}(x_{1},x_{2},y\mid m=0)}\log{}\left(\frac{p_{\Omega}^{\text{}}(x_{1},x_{2},y)}{p_{\Omega}^{\text{}}(x_{1},x_{2})p_{\Omega}^{\text{}}(y)}\right)\right]$
$=\mathbb{E}_{\begin{subarray}{c}x_{1},x_{2}\sim p_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2})\\ y\sim p_{\Omega^{\text{}}}(y\mid x_{1},x_{2})\end{subarray}}\left[\frac{p_{\Omega}^{\text{}}(x_{1},x_{2},y)}{\frac{p(m=0|x_{1},x_{2},y)p_{\Omega}^{\text{}}(x_{1},x_{2},y)}{p(m=0)}}\log{}\left(\frac{p_{\Omega}^{\text{}}(x_{1},x_{2},y)}{p_{\Omega}^{\text{}}(x_{1},x_{2})p_{\Omega^{\text{}}}(y)}\right)\right]$
$=\mathbb{E}_{\begin{subarray}{c}x_{1},x_{2}\sim p_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2})\\ y\sim p_{\Omega^{\text{}}}(y\mid x_{1},x_{2})\end{subarray}}\left[\frac{1-p(m=1)}{1-p(m=1|x_{1},x_{2},y)}\log{}\left(\frac{p_{\Omega^{\text{}}}(x_{1},x_{2},y)}{p_{\Omega^{\text{}}}(x_{1},x_{2})p_{\Omega^{\text{}}}(y)}\right)\right]$
That is, to estimate the mutual information under the true data distribution, we adjust for the shift in $p_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2})\mapsto p_{\Omega^{\text{}}}(x_{1},x_{2})$ and sample $y$ from the IPW-adjusted (parametrized approximations) of $p_{\Omega^{\text{}}}(y\mid x_{1},x_{2})$.
Appendix B Partial Information Decomposition (PID)
Partial information decomposition (PID Williams and Beer [48]) consists in decomposing the total mutual information [60, 61] between a target variable and two input variables into information about the target variable that both input variables share (“Shared” information), only one input variable has (“Unique” information) and emerges from the interactions of both (“Complementary” information). Bertschinger et al. [47] introduces bounds for these, reiterated below. In this Appendix, we express these bounds as entropy. First, Table 5 summarizes the notations used in the literature and those used in our work.
Table 5: Quantities and associated variables. Note that the four information measures are approximations.
Quantity Bertschinger ICYM2I
Input Variable 1 $Y$ $X_{1}$
Input Variable 2 $Z$ $X_{2}$
Target Variable $X$ $Y$
Redundant / Shared Information $\widetilde{SI}(X:Y:Z)$ $\widetilde{SI}(Y:X_{1};X_{2})$
[…]
Synergistic / Complementary Information $\widetilde{CI}(X:Y;Z)$ $\widetilde{CI}(Y:X_{1};X_{2})$
PID decomposition of the three-way mutual information $I(Y:(X_{1},X_{2}))$ results in the quantities of interest as follows:
$I(Y:(X_{1},X_{2}))=$ $\underbrace{SI(Y:X_{1};X_{2})}_{\text{Shared}}+\underbrace{UI(Y:X_{1}\backslash{X_{2}})}_{\text{Unique 1}}+\underbrace{UI(Y:X_{2}\backslash{X_{1}})}_{\text{Unique 2}}+\underbrace{CI(Y:X_{1};X_{2})}_{\text{Complementary}}$
Bertschinger et al. [47] provides the following bounds on each of these quantities:
$\widetilde{SI}(Y:X_{1};X_{2})$ $=\max_{q\in{\Delta_{\Omega^{\text{}}}}}CoI_{q}(Y;X_{1};X_{2})$
$=\max_{q\in{\Delta_{\Omega^{\text{}}}}}[I_{q}(Y:X_{1})-I_{q}(Y:X_{1}|X_{2})]$
$=\max_{q\in{\Delta_{\Omega^{\text{}}}}}\left[[I_{q}(Y:(X_{1},X_{2}))-I_{q}(Y:X_{2}|X_{1})]-I_{q}(Y:X_{1}|X_{2})\right]$
$=\max_{q\in{\Delta_{\Omega^{\text{}}}}}\left[I_{q}(Y:(X_{1},X_{2}))-[I_{q}(Y:X_{2}|X_{1})+I_{q}(Y:X_{1}|X_{2})]\right]$
[…]
$=\min_{q\in{\Delta_{\Omega^{\text{}}}}}[I_{q}(Y:(X_{1},X_{2}))-I_{q}(Y:X_{2})]$
[…]
$=\min_{q\in{\Delta_{\Omega^{\text{}}}}}[I_{q}(Y:(X_{1},X_{2}))-I_{q}(Y:X_{1})]$
$\widetilde{CI}(Y:X_{1};X_{2})$ $=I_{\Omega^{\text{}}}(Y:(X_{1},X_{2}))-\min_{q\in{\Delta_{\Omega^{\text{}}}}}I_{q}(Y:(X_{1},X_{2}))$
In this context, Bertschinger et al. [47] demonstrates that solving the optimization for $q\in{\Delta_{\Omega^{\text{}}}}$ that satisfies one of the four conditions above or the bound for conditional entropy, $\widetilde{H}$ formalized in (5), is sufficient to obtain all the quantities of interest.
$\widetilde{H}(Y|X_{1},X_{2})$ $=\max_{q\in{\Delta_{\Omega^{\text{}}}}}H_{q}(Y|X_{1},X_{2})$ (5)
Formulating PID quantities in terms of entropy
Remember that entropy $H(\cdot)$ is defined for general distributions of $X$ and $Y$ as follows:
$H(X)$ $:=-\sum_{x\in\mathcal{X}}p(x)\log{p(x)}$
$H(Y,X)$ $:=-\sum_{x\in\mathcal{X},y\in\mathcal{Y}}p(y,x)\log{\left(p(y,x)\right)}$
$H(Y|X)$ $:=-\sum_{x\in\mathcal{X},y\in\mathcal{Y}}p(y,x)\log{\left(\frac{p(y,x)}{p(x)}\right)}$
$=H(Y,X)-H(X)$
Using these notations, the mutual information $I(\cdot)$ can be defined as:
$I(Y:X)$ $:=H(X)-H(X|Y)$
$=H(X)+H(Y)-H(Y,X)$
$I(Y:X_{2}|X_{1})$ $:=H(Y,X_{1})+H(X_{1},X_{2})-H(Y,X_{1},X_{2})-H(X_{1})$
The previous quantities of interest can then be derived as:
$I(Y:(X_{1},X_{2}))$ $:=I(Y:X_{1})+I(Y:X_{2}|X_{1})$
$=\underbrace{H(Y)+\cancel{H(X_{1})}-\cancel{H(Y,X_{1})}}_{I(Y:X_{1})}+\underbrace{\cancel{H(Y,X_{1})}+H(X_{1},X_{2})-H(Y,X_{1},X_{2})-{\cancel{H(X_{1})}}}_{I(Y:X_{2}|X_{1})}$
$=H(Y)+H(X_{1},X_{2})-H(Y,X_{1},X_{2})$
where the first equation comes from the chain rule of mutual information [76].
Similarly, we can get the expression for co-information $CoI(Y;X_{1};X_{2})$:
$CoI(Y;X_{1};X_{2})=$ $I(Y:X_{1})+I(Y:X_{2})-I(Y:(X_{1},X_{2}))$
$=$ $[\underbrace{H(Y)-H(Y|X_{1})}_{I(Y:X_{1})}]+[\underbrace{H(Y)-H(Y|X_{2})}_{I(Y:X_{2})}]$
$\-[\underbrace{H(Y)+H(X_{1},X_{2})-H(Y,X_{1},X_{2})}_{I(Y:(X_{1},X_{2}))}]$
$=$ $[H(Y)-[H(Y,X_{1})-H(X_{1})]]+[\cancel{H(Y)}-[H(Y,X_{2})-H(X_{2})]]$
$-[\cancel{H(Y)}+H(X_{1},X_{2})-H(Y,X_{1},X_{2})]$
$=$ $H(Y)+H(X_{1})+H(X_{2})$
$-[H(X_{1},X_{2})+H(Y,X_{1})+H(Y,X_{2})]$
$+H(Y,X_{1},X_{2})$
The PID bounds can then be expressed in terms of entropy:
$\widetilde{SI}(Y:X_{1};X_{2})=$ $\max_{q\in{\Delta_{\Omega^{\text{}}}}}CoI_{q}(Y;X_{1};X_{2})$
$=$ $\max_{q\in{\Delta_{\Omega^{\text{}}}}}[H_{q}(Y)+H_{q}(X_{1})+H_{q}(X_{2})$
$-[H_{q}(X_{1},X_{2})+H_{q}(Y,X_{1})+H_{q}(Y,X_{2})]+H_{q}(Y,X_{1},X_{2})]$
[…]
$=$ $\min_{q\in{\Delta_{\Omega^{\text{}}}}}[H_{q}(Y,X_{2})+H_{q}(X_{1},X_{2})-H_{q}(Y,X_{1},X_{2})-H_{q}(X_{2})]$
[…]
$=$ $\min_{q\in{\Delta_{\Omega^{\text{}}}}}[H_{q}(Y,X_{1})+H_{q}(X_{1},X_{2})-H_{q}(Y,X_{1},X_{2})-H_{q}(X_{1})]$
$\widetilde{CI}(Y:X_{1};X_{2})=$ $I_{\Omega^{\text{}}}(Y:(X_{1},X_{2}))-\min_{q\in{\Delta_{\Omega^{\text{}}}}}I_{q}(Y:(X_{1},X_{2}))$
$=$ $[H_{\Omega^{\text{}}}(Y)+H_{\Omega^{\text{}}}(X_{1},X_{2})-H_{\Omega^{\text{}}}(Y,X_{1},X_{2})]$
$-\min_{q\in{\Delta_{\Omega^{\text{}}}}}[H_{q}(Y)+H_{q}(X_{1},X_{2})-H_{q}(Y,X_{1},X_{2})]$
where $H_{q}(\cdot{})$ and $H_{\Omega^{\text{}}}(\cdot{})$ is the entropy of a variable under probability distributions $q$ and $\Omega^{\text{}}$, respectively.
Appendix C ICYM2I-PID
Table 6: ICYM2I: Inverse probability weighting-adjusted multimodal training and evaluation under missingness shift.
Evaluation distribution
Observed Underlying
$\Omega_{\text{obs}}^{\text{}}$ $\Omega^{\text{}}$
[Uncaptioned image] [Uncaptioned image]
Training [Uncaptioned image]
Current
practice
IPW-adjusted
evaluation alone
IPW-adjusted
training alone
ICYM2I  (IPW-adjusted
training and evaluation)
Let $\text{IPW}_{q}(p)$ denote IPW reweighting to correct $p\mapsto q$ using samples from $p$. We use the following projection set to operationalize PID-bound estimation:
$\Delta_{\Omega^{\text{}}}^{\textsc{ICYM${}^{2}$I}}$
$\approx$ $\{q\propto\exp(f_{1}(x_{1})\cdot f_{2}(x_{2})):q(X_{i}=x_{i},Y=y)={p_{\Omega^{\text{}}}}_{\phi}(y,x_{i})\,\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}\}$
$=$ $\{q\propto\exp(f_{1}(x_{1})\cdot f_{2}(x_{2})):q(X_{i}=x_{i},Y=y)=\text{IPW}_{{p_{\Omega^{\text{}}}}_{\phi}}({p_{\Omega_{\text{obs}}^{\text{}}}}_{\phi}(y,x_{i}))\,\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}\}$
Note that this differs from prior works [46] that have used:
$\Delta_{\Omega^{\text{}}}\approx\{q\propto\exp(f_{1}(x_{1})\cdot f_{2}(x_{2})):q(x_{i},y)={\Omega_{\text{obs}}^{\text{}}}_{\phi}(x_{i},y)\,\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}\}$
To further improve the robustness of $p_{\Omega_{\text{obs}}^{\text{}}}(y,x_{i})\approx{p_{\Omega_{\text{obs}}^{\text{}}}}_{\phi}(y,x_{i})$, we optionally leverage both the unimodal and multimodal estimations, i.e., the unimodal and multimodal prediction models to enforce:
$\sum_{x_{\neg i}}f_{\phi}(y\mid x_{1},x_{2})p_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2})=\hat{f}_{\phi}(y\mid x_{i})p_{\Omega_{\text{obs}}^{\text{}}}(x_{i}),$
where $\hat{f}$ is obtained by adding the smallest amount of noise to $f_{\phi}(y\mid x_{i})\approx p_{\Omega_{\text{obs}}^{\text{}}}(y\mid x_{i})$ to ensure that the joint probabilities from the parametrized unimodal and multimodal models match exactly. This optimization did not improve PID values’ recovery.
We summarize our proposed auto-differentiable PID algorithm with IPW-based correction in Algorithm 1.
Algorithm 1 ICYM2I-PID
[…]
1:# Adjust $\Omega_{\text{obs}}^{\text{}}\mapsto\Omega^{\text{}}$.
2: Estimate missingness mechanisms ${p_{\Omega^{\text{}}}}_{\phi}(M_{1},M_{2},M_{Y}\mid C)$ for IPW.
3: Estimate corrected unimodal and multimodal models by training with weighting IPW-loss: $f_{\phi}(y\mid x_{i})\approx p_{\Omega^{\text{}}}(y\mid x_{i}),\,\forall i\in\{1,2\}$, and $f_{\phi}(y\mid x_{1},x_{2})\approx p_{\Omega^{\text{}}}(y\mid x_{1},x_{2})$.
4: # Drop subscript $\phi$ from $f_{\phi}$ for clarity.
5: Initialize parameterizations $\theta$ for $q$: $f_{i}(y\mid x_{i}),\,\forall i\in\{1,2\}$.
[…]
7:while not converged do
8:  for samples in batch do
9:    # Project onto $\Delta_{\Omega^{\text{}}}^{\textsc{ICYM${}^{2}$I}}$.
10:   $q_{\theta}(y\mid x_{1},x_{2})\leftarrow\texttt{SINKHORN-KNOPP}(q_{\theta}(y\mid x_{1},x_{2}),\{p(y,x_{i})\}_{i=1}^{2})$.
11:   Estimate the loss $I_{q}(Y:(X_{1},X_{2}))$ as a batch sample mean.
12:   $\theta\leftarrow\theta-\nabla_{\theta}I_{q}(Y:(X_{1},X_{2}))$.
13:  end for
14:end while# Estimate mutual information under $p_{\Omega^{\text{}}}$.
15: Estimate $I_{\Omega^{\text{}}}(Y:(X_{1},X_{2})),\,\text{and }I_{\Omega^{\text{}}}(Y:X_{i}),\,\forall i\in\{1,2\}$ using adjustment in Appendix A.
[…]
17:return $PID(\Omega^{\text{}})$
The traditional SINKHORN-KNOPP algorithm updates a matrix to enforce its marginals to be unit vectors. In our work, we adapt the algorithm to enforce the marginals to match $p_{\Omega}^{\text{}}$-marginals, ensuring that $q_{\theta}(\cdot)\in\Delta_{\Omega}^{\text{}}$. To ensure proper gradient propagation and reduce memory use, we use the unrolled SINKHORN-KNOPP [77, 78] algorithm. In the following, we use subscripts $q_{x_{1},x_{2}}$ to denote $q_{\theta}(y,x_{1},x_{2})$ and $p_{x_{i}}$ to denote $p_{\phi}(y,x_{i})$. The algorithm is detailed below:
Algorithm 2 Unrolled SINKHORN-KNOPP update
[…]
3:while do
4:   # Avoid update if both exit conditions have been met.
5:  if $\left|\frac{q_{x_{1}}-p_{x_{1}}}{p_{x_{1}}}\right|\leq\texttt{atol}$ $\left|\frac{q_{x_{2}}-p_{x_{2}}}{p_{x_{2}}}\right|\leq\texttt{atol}$ then
6:   return $q_{x_{1}x_{2}}$
7:  end if
8:   # Update marginal.
9:  $q_{x_{1}x_{2}}\leftarrow\frac{q_{x_{1}x_{2}}}{q_{x_{2}}}\cdot p_{x_{2}}$
10:  $q_{x_{1}}\leftarrow\sum_{x_{2}}q_{x_{1}x_{2}}$
11:   # If the other marginal still matches, done.
12:  if $\left|\frac{q_{x_{1}}-p_{x_{1}}}{p_{x_{1}}}\right|\leq\texttt{atol}$ then
13:   return $q_{x_{1}x_{2}}$
14:  end if
15:   # Repeat for the other marginal.
16:  $q_{x_{1}x_{2}}\leftarrow\frac{q_{x_{1}x_{2}}}{q_{x_{1}}}\cdot p_{x_{1}}$
17:  $q_{x_{2}}\leftarrow\sum_{x_{1}}q_{x_{1}x_{2}}$
18:  if $\left|\frac{q_{x_{2}}-p_{x_{2}}}{p_{x_{2}}}\right|\leq\texttt{atol}$ then
19:   return $q_{x_{1}x_{2}}$
20:  end if
21:end while
Appendix D Bit-wise logits
In this section, we perform a sensitivity analysis of the logit setting presented in Section 4.1 under two additional missingness patterns: MCAR (Missing Completely at Random) and MNAR (Missing Not at Random). In this setting, we fit a logistic regression to estimate the probability of missingness on the observed modality, which is then used to estimate the IPW. The performance estimates and PID for these two missingness processes are illustrated in Tables 7 and 8.
Since MCAR does not result in a distribution shift, one expects the same performance estimates for both the full and observed populations. Furthermore, in this setting, the IPW correction corresponds to a constant value, as any point has the same probability of observing both modalities. This correction also results in no change in performance estimates.
On the contrary, MNAR patterns do not guarantee similar behavior. Particularly, this missingness process may result in a distribution shift that cannot be assessed or accounted for without assumptions about the data distribution, as one does not observe the covariates that impact the missingness process. The results demonstrate that both the observed and corrected strategies result in biased estimates.
Table 7: Impact of missingness on multimodality information for bitwise logic operators under MCAR. Parentheses denote standard deviation across batches.
AUROC Information Decomposition
$X_{1}$ $X_{2}$ $X_{1}+X_{2}$ Unique 1 Unique 2 Shared Complementary
AND Oracle 0.83 (0.01) 0.84 (0.01) 1.00 (0.00) 0.05 (0.00) 0.03 (0.00) 0.26 (0.00) 0.47 (0.00)
Observed 0.83 (0.01) 0.83 (0.01) 1.00 (0.00) 0.05 (0.00) 0.03 (0.00) 0.23 (0.00) 0.52 (0.00)
ICYM2I 0.83 (0.01) 0.85 (0.01) 1.00 (0.00) 0.03 (0.00) 0.06 (0.00) 0.27 (0.00) 0.44 (0.00)
OR Oracle 0.84 (0.01) 0.83 (0.01) 1.00 (0.00) 0.04 (0.00) 0.05 (0.00) 0.27 (0.00) 0.46 (0.00)
Observed 0.84 (0.01) 0.84 (0.01) 1.00 (0.00) 0.06 (0.00) 0.03 (0.00) 0.25 (0.00) 0.51 (0.00)
ICYM2I 0.85 (0.01) 0.83 (0.01) 1.00 (0.00) 0.06 (0.00) 0.02 (0.00) 0.25 (0.00) 0.51 (0.00)
XOR Oracle 0.51 (0.02) 0.49 (0.01) 1.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.99 (0.00)
Observed 0.51 (0.02) 0.50 (0.02) 1.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.95 (0.00)
ICYM2I 0.51 (0.02) 0.51 (0.02) 1.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.95 (0.00)
Table 8: Impact of missingness on multimodality information for bitwise logic operators under MNAR. Parentheses denote standard deviation across batches.
AUROC Information Decomposition
$X_{1}$ $X_{2}$ $X_{1}+X_{2}$ Unique 1 Unique 2 Shared Complementary
AND Oracle 0.83 (0.01) 0.84 (0.01) 1.00 (0.00) 0.05 (0.00) 0.03 (0.00) 0.26 (0.00) 0.47 (0.00)
Observed 0.93 (0.01) 0.67 (0.01) 1.00 (0.00) 0.45 (0.00) 0.00 (0.00) 0.17 (0.00) 0.33 (0.00)
ICYM2I 0.93 (0.01) 0.67 (0.01) 1.00 (0.00) 0.45 (0.00) 0.00 (0.00) 0.17 (0.00) 0.33 (0.00)
OR Oracle 0.84 (0.01) 0.83 (0.01) 1.00 (0.00) 0.04 (0.00) 0.05 (0.00) 0.27 (0.00) 0.46 (0.00)
Observed 0.78 (0.01) 0.95 (0.01) 1.00 (0.00) 0.00 (0.00) 0.17 (0.00) 0.11 (0.00) 0.23 (0.00)
ICYM2I 0.78 (0.01) 0.95 (0.01) 1.00 (0.00) 0.00 (0.00) 0.17 (0.00) 0.11 (0.00) 0.23 (0.00)
XOR Oracle 0.51 (0.02) 0.49 (0.01) 1.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.00 (0.00) 0.99 (0.00)
Observed 0.80 (0.02) 0.52 (0.02) 1.00 (0.00) 0.35 (0.00) 0.07 (0.00) 0.00 (0.00) 0.61 (0.00)
ICYM2I 0.80 (0.02) 0.52 (0.02) 1.00 (0.00) 0.35 (0.00) 0.07 (0.00) 0.00 (0.00) 0.61 (0.00)
Appendix E Synthetic data results
Data generation. Our work builds on the example introduced in [46], in which we enforce additional missingness. Three latent variables ($z_{1}$, $z_{2}$, and $z_{c}$) are drawn from multi-dimensional clustered data; the observed covariates are a concatenation of $z_{c}$ and one of the other latent variables, as illustrated in Figure 3. Then, the outcome $Y$ is generated as $Y=\sigma(p_{1}\mathbb{E}(z_{1})+p_{2}\mathbb{E}(z_{2})+(1-p_{1}-p_{2})\mathbb{E}(z_{c}))$, with the proportion $p_{i}\in[0,1]$ such that $p_{1}+p_{2}\leq 1$. We simulate datasets with varying values of $p_{1}$ and $p_{2}$. Then, we enforce a $50\%$ MAR missingness pattern in $X_{2}$ by modeling the probability of missingness. We do this by clustering $X_{1}$ into 100 groups using Kmeans. Then, the probability of missingness is generated using a random forest that regresses $X_{1}$ to predict $c_{j}\cdot{Y}$.
$z_{c}$$z_{2}$$z_{1}$$X_{1}$$X_{2}$$Y$
Figure 3: Data generating processes for synthetic experiments. $z_{i}$ denote latent vectors, while all other variables are observed. Filled point nodes are observed variables, while unfilled nodes are unobserved.
Empirical setting. Data were split into three: 80% for training, 10% for validation, and the rest for testing. We consider neural networks with 2 hidden layers with 32 nodes, trained using an Adam optimizer [79] with a learning rate of 0.001 over 100 epochs. Our evaluation relies on discriminative performance measured through AUROC.
Estimating predictive performance under $\Omega_{\text{obs}}^{\text{}}$. Table 9 presents the estimated performance obtained under different corrections. These results underline the importance of correcting both training and evaluation, as proposed in ICYM2I, to best align with the performance one would obtain on $\Omega^{\text{}}$, as shown by the smallest Root Mean Squared Error (RMSE) observed when both corrections are applied. Note that in this setting, we rely on the true IPW correction that one would obtain with a properly specified model, as the MAR setting is met.
Table 9: Comparison between estimated AUC performance under the different training and evaluation corrections and oracle performance on $\Omega^{\text{}}$. $\epsilon$ denotes the RMSE between estimated and oracle PIDs.
Evaluation
Observed Underlying
$\Omega_{\text{obs}}^{\text{}}$ $\Omega^{\text{}}$
[Uncaptioned image] [Uncaptioned image]
Training [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Appendix F Semi-synthetic data results
F.1 UR-FUNNY
We illustrate the impact of missingness on estimating the informativeness of different modalities on real-world data with UR-FUNNY [64], a multimodal dataset for humor detection from human speech used in affective computing. The dataset comprises text, audio, and visual modalities from 10 - 20 second videos sourced from TED talks, and the task is to detect whether a punchline would trigger a laugh. Labels were generated using the markup “(Laughter)” [80] from the transcript.
Dataset. The processed dataset from MultiBench [81] is a modality-complete dataset with 10,166 samples of paired audio, text, and vision embeddings. Audio embeddings were generated with COVAREP [82], text with Glove [83], and visual features through the Facet [84] library and OpenFace [85], and aligned using the Penn Phonetics Lab Forced Aligner (P2FA) [84].
Enforcing missingness. To explore the impact of missingness on informativeness, we simulate a MAR missingness pattern on the audio and visual features given the textual modality. We vary the missingness from 30% to 70%, using the same mechanism as described for synthetic data. This semi-synthetic setting enables the evaluation of the proposed correction as the missingness mechanism is known. Note that the original dataset does not contain missing values, as the source data (TED Talks) have transcripts, and data labeling was generated based on these transcripts. However, settings with systematic transcripts are rare and may reflect a shift from the audio and textual modalities observed online for which such a match may not exist.
Results. Following the same empirical setting as in the synthetic experiment for each missingness rate, we measure the impact of missingness on PID decomposition. Figure 4 displays the PID values obtained under three strategies:
  • Observed: All quantities are estimated using $\Omega_{\text{obs}}^{\text{}}$.
  • ICYM2I: All quantities are estimated using $\Omega_{\text{obs}}^{\text{}}$ but corrected for the distribution shift through IPW.
  • Oracle: All quantities are estimated on $\Omega^{\text{}}$.
This figure shows that the proposed strategy is consistently closer to the Oracle’s PID values. This demonstrates that under Assumption C, the proposed correction yields better estimates of each modality’s informativeness – specifically, the audio-visual modality (Unique 1) carries more information.
Figure 4: Comparison between estimated PID values under increasing missingness in UR-FUNNY.
F.2 Hateful Memes
We run experiments using the dataset from the Hateful Memes Challenge [65], which investigates text-image multimodal reasoning in the context of hate speech detection in online memes. The dataset comprises text-image pairs with an associated label indicating hate speech.
Dataset. We utilize the Kaggle version of the Facebook Hateful Memes dataset, as referenced in the Holistic Evaluation of Multimodal Foundation Models (HEMM) [86] repository. Our analysis focuses on the 9,000 samples with associated labels. For each sample, embeddings were extracted for both modalities using a ResNet-50 [87] for images and a BERT-base-uncased [88] model for text. The proposed ResNet-50 was pretrained on ImageNet [89] with the final layer replaced to extract 2048-dimensional feature vectors, and BERT-base-uncased [88] extracts embeddings of dimension 784 from the penultimate layer.
Enforcing missingness. Similarly to the previous experiment, we vary the missingness from 30% to 70% by enforcing the same MAR missingness mechanism on the text modality, given the image modality, as we assume not all memes may contain text. Note that memes in the dataset were created by combining text from collected online memes with images sourced from stock images on Getty Images. Consequently, the dataset did not contain missing modality, but may not match the true distribution of memes one would observe online.
Results. As above, we measure the impact of increasing percentages of missingness on PID estimates. While the missingness mechanism results in a limited distribution shift, and therefore small differences in estimates between the corrected and observed strategies, the difference at 70% missingness shows the superiority of the proposed methodology in recovering the Unique contributions.
Figure 5: Comparison between estimated PID values under increasing missingness in Hateful Memes.
Appendix G Structural heart disease data processing
All electrocardiograms were 10-second, standard 12-lead ECG signals collected at abstracted to 250 Hz, which we resampled to 500 Hz, and standard normalized by channel to match the inputs for ECG-FM [71]. We used the version of ECG-FM with weights pretrained on MIMIC-IV [90, 91] and PhysioNet 2021 [92, 93]. We averaged the outputted feature embeddings along the temporal dimension and flattened them to produce vectors of length 768.
The chest radiographs used in our study were all postero-anterior (PA) view CXRs. We extracted pixel values from the DICOM files as grayscale images, center-cropped each image along the shorter dimension, applied contrast-limited adaptive histogram equalization (CLAHE) [94] with a clip limit of 0.2, and resized each image to $1284\times{}1284$ pixels. All outputted embeddings were flattened to 4098-dimensional vectors.
Appendix H Compute infrastructure
All experiments were performed on a server with an AMD EPYC 7313 CPU, 256 GB of memory, and two NVIDIA RTX A6000 GPUs, as well as a server with an Intel Xeon E5-2640 CPU, 128 GB of memory, and a NVIDIA GTX Titan X GPU. Our software stack includes Python 3.12, PyTorch 2.2.1 [95], and standard Python scientific libraries. Chest radiograph embeddings used Tensorflow 2.19 [96] and Tensorflow-Text 2.19 based on the requirements for the ELIXR models [72]. Electrocardiogram embeddings were generated using an environment with Python 3.9 and fairseq-signals 1.0 to match the requirements for fairseq-signals and ECG-FM [71]. Generating embeddings for our structural heart disease data took approximately 10 hours on our server with a Titan X GPU. All synthetic experiments require 12 hours of compute time using one NVIDIA RTX A6000 GPU.